OpenStudy (anonymous):

can someone show me by drawing what csc^-1 sqrt 2=?

OpenStudy (anonymous):

@dan815 @Nnesha @mathmate

OpenStudy (anonymous):

inverse trigonometry... anyone?

OpenStudy (anonymous):

arccsc or csc - 1?

OpenStudy (anonymous):

arc csc

OpenStudy (anonymous):

Are you given anything else? arcsc is simply \(\sf \large \frac{1}{arcsine}\)

OpenStudy (anonymous):

oups: \[\csc ^{-1}\sqrt{2}\]

OpenStudy (loser66):

it is just pi/4 , 3pi/4, 5 pi/4 and 7pi/4

OpenStudy (loser66):

\(csc t = \sqrt 2\\csc t = \dfrac{1}{sin t}\) and we know that \(csc t = \sqrt 2\implies arccsc \sqrt 2= t\)

OpenStudy (loser66):

Now, back to csc \(csc t =\dfrac{1}{sin t}=\sqrt 2\), then s\(int = \sqrt2/2 \) that is t = pi/4, 3pi/4.....\)

OpenStudy (loser66):

Got me?

OpenStudy (loser66):

Yes or no? need more explanation?

OpenStudy (anonymous):

wait up... trying to fully understand here...

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[Let~ \csc ^{-1}\sqrt{2}=y,\csc y=\sqrt{2},\sin y=\frac{ 1 }{ \sqrt{2} }=\sin \frac{ \pi }{ 4 },\sin \left( \pi-\frac{ \pi }{ 4 } \right)\] y=?